OpenStudy (ell):

Use the properties of exponents to simplify each expression. 2(3x)^3

OpenStudy (jim766):

2(3^3)(x^3) now simplify

OpenStudy (ell):

2*3=6 2*x=2x 2+3=6 2+3=6 ?

OpenStudy (jim766):

need to follow order of operations ( )'s first 2 (3^3)(x^3) 2 (27)(x^3) 52x^3 make sense?

OpenStudy (ell):

Yeah! haha(: thanx for taking the time to help! ur awesome! Much appreciated! GBU!! <3 :)